Let h(n) =1 if n codes a Turing machine M which halts when started on a blank tape, h(n) =0 otherwise. Sketch a proof that h is not Turing computable.
Answer #1

h is not turing compatable.why means everytime we will get the turing resukt as 1 only

if n=1 then turing give resulkt as h(1)=1 then turing give resukt as 1=1 o/p as 11=11

if n=2 then turing give resulkt as h(2)=1 then turing give resukt as 2=1 o/p as 111=11

if n=3 then turing give resulkt as h(3)=1 then turing give resukt as 3=1 o/p as 1111=11

if n=0 then turing give resulkt as h(0)=1 then turing give resukt as 0=1 o/p as 1=11

so evertime we will get output as 11 only at right side.so h(n)=0 is not turing compatable